# Digging into contrastive learning for robust depth estimation with diffusion models

### v3

@@ 4.4. Ablation Study
In this section, we conduct detailed ablation studies on WeatherKITTI and real-world benchmarks to demonstrate the effectiveness of the proposed components.
Table 4. Abladation study on stable depth with diffusion. All experiments are implemented in stage one and the meaning of the abbreviation is as follows. PDE: Pseudo-depth distillation enhancement; ODR: Outlier depth removing; FIC: Feature-image joint condition.
@@
Table 5. Abladation study on contrast manner and level. All experiments are implemented in stage second and the meaning of the abbreviation is as follows. NIS: the noise level contrast; DF: the depth feature level contrast; IMG: the image level contrast. Real Weather shows the average test result across all kinds of real weather conditions in Tab.2
@@
Effectivness of Diffusion Enhancement. Table 4 shows the effectiveness of each enhancement based on the first stage of training. We take the MonoDiffusion (replaced with MonoViT) as the baseline (ID 1). Firstly, the distillation enhancement (ID 2) is added and produces a significant improvement. This gain is reasonable because, compared with MonoDiffusion which only considers clear scenes, the depth estimation in complex scenes typically requires more powerful distillation. Next, we add outlier depth removal (ID 3), which brings a notable improvement in RMSE and SqRel metrics because the unnatural depth values have more significant impact on the second-order indicators. Finally, we improved the condition (ID 4). With the introduction of image context information, the diffusion-based model further shows an improvement in the $a_{1}$ metric representing estimation stability.
Effectiveness of Trinity Contrast. As reported in Table 4 (ID 5), the incorporation of trinity in the first stage model produces overall performance gains in all aspects. Meanwhile, in Table 5, we can compare the performance of three paradigms using IDs 4-6 and IDs 10-12. The results show that direct distillation did not enhance the performance and even led to a decrease in stability (RMSE, $a_{1}$). For the contrastive method, although we applied the constraint of Eq. 2 to the contrastive objects, directly contrasting the predicted noises would lead to network crashes (i.e., in which all predicted depth values are zeros, which are not included in Table 5). Therefore, we still maintain the noise constraint from Eq. 6, only taking the feature and image levels comparison, but there is only a minor improvement. The proposed trinity contrast is the only approach that significantly improves convergence on WeatherKITTI and demonstrates increased robustness and accuracy on real datasets.
Effectiveness of Multi-Level Contrasts The comparison of ablation with multi-level contrasts is presented in Table 5 (IDs 2,3,6 and IDs 8,9,12). While the performance of different models is close in the WeatherKITTI dataset, in real-world weather situations, we notice that the results become more reliable as we gradually include the feature and image level trinity. The enhancements in these metrics, coupled with Fig. 4, apparently prove the effectiveness of the multi-level contrast.
@@ 5. Conclusion
In this paper, we enhance the stability and convergence of diffusion-based MDE and further introduce a novel robust depth estimation framework named D4RD, which integrates a customized contrastive learning scheme. This method combines the strength of distillation with contrastive learning and repurposes the sampled noise in the diffusion forward process, creating a ”trinity” contrastive mode, thereby improving the accuracy and robustness of noise prediction. Furthermore, we extend the noise level’s trinity to the more general feature and image levels, constructing a multi-level trinity scheme to balance the robust perception pressure across different model components. Through extensive quantitative and qualitative experiments, we demonstrate the effectiveness of D4RD against various architectures and its superior performance over existing SoTA solutions.

### v4

@@ 4.4. Ablation Study
In this section, we conduct detailed ablation studies on WeatherKITTI and real-world benchmarks to demonstrate the effectiveness of the proposed components. In the end, we discuss the computational analysis and training costs of D4RD.
Table 5. Ablation study on stable depth with diffusion. All experiments are implemented in stage one, and the meaning of the abbreviations is as follows: PDE stands for Pseudo-depth Distillation Enhancement; ODR stands for Outlier Depth Removing; FIC stands for Feature-Image joint Condition.
@@
Table 6. Ablation study on contrast manner and level. All experiments are implemented in the second stage and the meaning of the abbreviations is as follows. NIS: the Noise level contrast; DF: the Depth Feature level contrast; IMG: the Image level contrast. Real Weather shows the average test result across all kinds of real weather conditions in Tab.2
@@
Effectivness of Diffusion Enhancement. Table 5 shows the effectiveness of each enhancement based on the first stage of training. We take the MonoDiffusion (replaced with MonoViT) as the baseline (ID 1). Firstly, the distillation enhancement (ID 2) is added and produces a significant improvement. This gain is reasonable because, compared with MonoDiffusion which only considers clear scenes, the depth estimation in complex scenes typically requires more powerful distillation. Next, we add outlier depth removal (ID 3), which brings a notable improvement in RMSE and SqRel metrics because the unnatural depth values have a more significant impact on the second-order indicators. Finally, we improve the condition (ID 4). With the introduction of image context information, the diffusion-based model further shows an improvement in the $a_{1}$ metric representing estimation stability.
Effectiveness of Trinity Contrast. As reported in Table 5 (ID 5), the incorporation of trinity in the first stage model produces overall performance gains in all aspects. Meanwhile, in Table 6, we can compare the performance of three paradigms using IDs 4-6 and IDs 10-12. The results show that direct distillation did not enhance the performance and even led to a decrease in stability (RMSE, $a_{1}$). For the contrastive method, although we applied the constraint of Eq. 2 to the contrastive objects, directly contrasting the predicted noises would lead to network crashes (i.e., all predicted depth values are zeros, which are not included in Table 6). Therefore, we still maintain the noise constraint from Eq. 6, only taking the feature and image levels comparison, but there is only a minor improvement. The proposed trinity contrast is the only approach that significantly improves convergence on WeatherKITTI and demonstrates increased robustness and accuracy on real datasets.
Effectiveness of Multi-Level Contrasts. The comparison of ablation with multi-level contrasts is presented in Table 6 (IDs 2,3,6 and IDs 8,9,12). While the performance of different models is close in the WeatherKITTI dataset, in real-world weather situations, we notice that the results become more reliable as we gradually include the feature and image level trinity. The enhancements in these metrics, coupled with Fig. 4, apparently prove the effectiveness of the multi-level contrast.
Compute Analysis. As shown in Table 4, D4RD’s diffusion process increases parameter count by 0.4% and computational effort by 23.3% compared to the baseline (traditional MDE). However, this is justified by a 17.5% reduction in AbsRel error. The training process does incur higher memory and time expenses, but these can be mitigated by using multiple GPUs, and training is a one-time process. For inference, D4RD is slower due to iterative denoising, but several acceleration techniques have been successfully applied to solve these(see (Ke et al., 2024)’s GitHub). Additionally, the two-stage strategy is employed only during training, whereas the inference process remains end-to-end.
@@ 5. Conclusion
In this paper, we enhance the stability and convergence of diffusion-based MDE and further introduce a novel robust depth estimation framework named D4RD, which integrates a customized contrastive learning scheme. This method combines the strengths of distillation with contrastive learning and repurposes the sampled noise in the diffusion forward process, creating a ”trinity” contrastive mode, thereby improving the accuracy and robustness of noise prediction. Furthermore, we extend the noise level’s trinity to the more general feature and image levels, constructing a multi-level trinity scheme to balance the robust perception pressure across different model components. Through extensive quantitative and qualitative experiments, we demonstrate the effectiveness of D4RD against various architectures and its superior performance over existing SoTA solutions.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Nos. 62172032).
Supplementary Materials
6. Evaluation Metrics
Similar to (Zhao et al., 2022b), we employ the following evaluation metrics in our experiments,
  • AbsRel: $\ \frac{1}{|\mathcal{M}|}\sum_{d\in\mathcal{M}}\left|d-d_{gt}\right|/d_{gt}$;
  • SqRel: $\frac{1}{|\mathcal{M}|}\sum_{d\in\mathcal{M}}\left\|d-d_{gt}\right\|^{2}/d_{gt}$;
  • RMSE: $\sqrt{\frac{1}{|\mathcal{M}|}\sum_{d\in\mathcal{M}}\left\|d-d_{gt}\right\|^{2}}$;
  • RMSElog: $\sqrt{\frac{1}{|\mathcal{M}|}\sum_{d\in\mathcal{M}}\left\|\log(d)-\log(d_{gt})\right\|^{2}}$;
  • $\textbf{a}_{t}$: percentage of $d$ such that $\max(\frac{d}{d_{gt}},\frac{d_{gt}}{d})<1.25^{t}$
where $d_{gt}$ and $d$ denote the GT and estimated pixel depth, $\mathcal{M}$ is the valid mask set to $801e-3<d_{gt}<$ 80. Note that we also use the median scaling technique introduced by (Zhou et al., 2017) on $d$ to recover the absolute depth scale.
7. Formula symbol summary
In this paper, to explain our method clearly, we use numerous symbols in the formulas and the meaning of each symbol can be found at its first appearance. However, for the readers’ convenience, we once again list the symbol meanings in Table. 7. Note that we only record meta symbols here and do not include compound symbols with superscripts and subscripts. The meaning of any symbol in this paper can be grasped through combinations. (e.g. $\textbf{d}_{S,clr}$ denote the depth prediction of the clear images from the student model).
Table 7. The meanings of meta symbols. Because the same character can represent multiple meanings, please pay attention to font changes.
Symbol Meaning
a The auxiliary image from adjacent frame
$\alpha$ A transformation of $\beta$
aug The augmented image
$\beta$ Predefined variance noise schedule
c The diffusion’s condition
clr The clear image
cst Contrast (a learning method)
d Pixel level depth
d Map level depth prediction
D Map level depth distribution (from GT)
dis Distillation (a learning method)
e The edges
$\epsilon$ A noise that follows $\mathcal{N}$(0, I) distribution
f, feat The depth feature
$\mathcal{F}$ A neural network, the concrete meaning is
determined by its index
gt The ground truth depth
I, img The RGB images
L The training loss
M Adaptive distillation mask
$\mathcal{M}$ The valid mask of GT depth
$\mathcal{N}$ Gaussian sampling
nis The noise
ph Photometric consistency
S Student model
T Teacher model
$\mathcal{T}$ Relative camera poses transformation
$\mathbb{T}$ Overall diffusion training steps
t The target image (The image to be estimated depth)
$\tau$ The time step
$\mathcal{U}$ Uniformly sampling
$\lambda$ A constant, relative to adaptive mask
$\omega,\eta,\rho$ Weights for various losses
8. The snowy dataset
Figure 7. The LiDAR points projection on the CADC dataset and Dense dataset. The cropped areas and prominent errors are highlighted with green and orange boxes, respectively. For the CADC dataset, the colors from near to far are red, orange, yellow, green, blue, and purple in order. And for Dense, it is blue, light blue, yellow, orange, and red.
In WeatherDepth(Wang et al., 2024), Wang et al. use the CADC dataset to validate their model’s performance in snowy scenes. However, as shown in Fig. 7(a), the LiDAR point clouds in this dataset have many errors on pedestrians and road signs. Furthermore, there is a vehicle bonnet at the bottom of each image, so Wang et al. have to crop them and the sky part, which greatly disrupts the depth clues brought by the vertical position. Hence, the experiments on this dataset can only give a rough result and we need a better depth-annotated snowy dataset.
In this work, we chose the Dense dataset. As depicted in Fig. 7(b), the LiDAR points in Dense are more accurate and only require cropping a small portion of the image. All of these factors make the comparison result reasonable.
9. Additional Qualitative Results
Figure 8. Extra visual comparisons on synthetic and real weather dataset. It is better viewed when zooming in. (a)-(g) are the WeatherKITTI dataset results and (i)-(m) are the real weathers.
To further vividly demonstrate the superiority of our method, as shown in Fig. 8, we present a visual comparison of each scene from both synthetic and real datasets. (a) is clear scene. The highlighted area in the dashed box exemplifies how pedestrians wearing white clothing are accurately distinguished from the background by D4RD. (b) is synthetic light rainy scene. Despite being affected by camera blur, D4RD still recognizes the correct shape of the streetlights. (c) is synthetic light snowy scene. D4RD stands out as the only RMDE model capable of not mistaking train windows for background depth. (d) is synthetic light foggy scene. D4RD correctly recognizes that the wall is a plane, avoiding concave depth errors. (e) is synthetic adverse rainy scene. The influence of railway tracks on depth should be negligible, but under the interference of raindrop particles, only D4RD exhibits correct depth structures. (f) is synthetic adverse snowy scene. D4RD not only identifies the white car far away but also recognizes the almost invisible road sign (on the left side of the image). (g) is synthetic adverse foggy scene. From the clear image, it can be seen that the object highlighted by the green box is a distant building roof, which should not be recorded in the depth map (given a maximum depth of 80m). Only D4RD can discern such details accurately. (i) is the real rainy scene and a failure case for all RMDE models. Weatherdepth and Robust-Depth* identified lens blurring as infinity, while EC-Depth* did not recognize the sparse areas. Although D4RD addresses these issues to some extent, it remains flawed. This imperfection is likely attributed to insufficient training data available for this specific setting. (j) is real snowy scene. Once again, the distant cars and fences are accurately identified by D4RD. (k) is real cloudy scene. From the zoomed image, it can be seen that the building should be located behind a distant pole, which exceeds the range of 80m. However, D4RD can accurately identify this case. (l) is from a real sunny scene, specifically focusing on the unique condition of the bridge cave. Affected by this out-of-distribution environment, the depth estimation performance of other RMDE models was significantly impacted. Conversely, D4RD demonstrates robustness and maintains accurate depth estimation. (m) is real fog scene. From the zoomed image, it can be seen that the correct structural relationship of the highlighted area is ‘a pole in front of the distant building’. Other models either estimate the depth of the building too closely or estimate the depth of the poles too far. Only D4RD is competent for this job.
Based on the above discussion, we further prove the robustness and effectiveness of D4RD.